OpenStudy (anonymous):

the area of a rectangle is 60 in.2 and the width is 4 in. what is the rectangle's length? can some one help me understand this?

Mehek (mehek14):

4 * 2?

OpenStudy (anonymous):

8

Mehek (mehek14):

do you have any answer choices first of all?

OpenStudy (anonymous):

15 in, 26 in 52 in

OpenStudy (phi):

do you know the formula for the area of a rectangle is L*W = Area where L is the length and W is the width ?

OpenStudy (anonymous):

the formula my teacher gave me for rectangle is: p=2l+2w and a=lw

OpenStudy (phi):

they say width is 4 in so W is 4 and area of a rectangle is 60 so the area is 60 if we use the formula W*L= area 4 * L = 60 that says 4 times a number is 60 (the number is the length) which choice for L makes 4*L be 60 ?

OpenStudy (phi):

if the width is 4 and the length is 15, what area do you get ? can you do that ?

OpenStudy (anonymous):

i got 15

OpenStudy (phi):

yes if the length is 15 and the width is 4, then the area is 4*15= 60 which matches the problem. this means the length is 15